A basketball is shot through the air, hitting the backboard of a hoop. During the shot, the basketball was travelling at 8 m/s, and right after hitting the backboard, the basketball travels 1 m/s in the opposite direction. The mass of a standard basketball is 590g.



a. Find the kinetic energy of the basketball before hitting the backboard (2pts)

b. Find the kinetic energy of the basketball after hitting the backboard (2pts)


1
Expert's answer
2021-11-30T17:43:56-0500

a. "KE_1=mv_1^2/2=0.59\cdot8^2/2=18.88\ (J)"


b. "KE_2=mv_2^2/2=0.59\cdot1^2/2=0.295\ (J)"